OpenStudy (anonymous):

A trapezoid has an area of 24 square meters. The lengths of the bases of the trapezoid are 5 meters and 7 meters. What is the height of the trapezoid?

Directrix (directrix):

The area of a trapezoid is: A = (1/2)*h* (sum of the two bases) 24 = (1/2) * h * ( 5 + 7) Multiply both sides by 2. 48 = 12* h Solve for h. h = ?